OpenStudy (anonymous):

Use basic identities to simplify the expression. (cos^2(theta))/(sin^2(theta))+ csc(theta) sin(theta)

OpenStudy (anonymous):

\[\frac{ \cos^2(\theta) }{ \sin^2(\theta) }+ \csc(\theta) \sin(\theta)\]

OpenStudy (dan815):

duudeee

OpenStudy (dan815):

csc=1/sin

OpenStudy (dan815):

cos^2/sin^2 = 1/tan^2 or cot^2

OpenStudy (dan815):